[figure: 9.7 cm]

ANOMALOUS HALL EFFECT IN MAGNETIC KONDO-LATTICE CeAl_{2}

N.E.Sluchanko^{a}, A.V.Bogach^{a}, V.V.Glushkov^{a}, S.V.Demishev^{a}, G.S.Burkhanov^{b}, O.D.Chystyakov^{b} aGeneral Physics Institute of RAS, Vavilov str., 38, Moscow, 119991, Russia

bInstitute of Metallurgy and Material Science of RAS, Leninskii Prospect, 49, Moscow, 117334, Russia

A wide temperature range (1.8-300K) Hall coefficient (R_{H}) measurements have been carried out in magnetic fields up to 8T on a high quality polycrystalline samples of a magnetic Kondo-lattice CeAl_{2}. The investigation of angle and field dependencies of the Hall coefficient and magnetoresistance in paramagnetic and modulated antiferromagnetic phases of CeAl_{2} allows to deduce a very complicated Hall effect consisting of normal and several anomalous contributions (spin- fluctuation and magnetic components). Among them the anomalous positive component of R_{H} (spin-fluctuation contribution) demonstrates a broad maximum around T≈4K which is depressed dramatically (by a factor of 3) in magnetic field H≈8T. The anomalous magnetic components of the Hall coefficient R_{H}^{m} are characterized by a narrow maximum at T=T_{N}≈3.85K and a nonmonotonous behavior of R_{H}^{m} in a magnetic field below 4T which likely may be attributed to the AFM-domains reorientation process at liquid helium temperatures.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : A.V.Bogach

Address for correspondence : General Physics Institute of RAS, Vavilov str., 38, Moscow, 119991, Russia

Email address : nes@lt.gpi.ru